You are experimenting with an ethanol and water mi
Answer #1

a)
delta H vap can calculated from graph and also by formula:
y = -3.7052 x + 42.516
= -3.7052 *0.61 + 42.516
=40.25

Also from graph draw a line and 0.61 and then a horizintal line to get 40.25 KJ/mol

Answer: 40.25 KJ/mol

b)
P1=1.04 atm
T1 = (273 + 23.6) K = 296.6 K
P2= ?
T2 = (273+34.4) K = 307.4 K

use:
P1/T2 = P2/T2
1.04/296.6 = P2/307.4
P2 = 1.08 atm

Answer: 1.08 atm